OpenStudy (deez):

Show that any separable equation M(x) + N(y) yَ = 0 is also exact

OpenStudy (anonymous):

Not any separable equation is exact.. well let's see: M(x) + N(y) (dy/dx) = 0 Multiply across by dx M(x)dx+N(y)dy=0.

OpenStudy (anonymous):

that equation is going to be exact if and only if: Mx = Ny

OpenStudy (deez):

then how im gonna proof that ??? its a problem in Willian E. Boyce's Book Elementary Differential Equations \ 7th Edition

OpenStudy (anonymous):

No just take the function F = int M(x)dx + int N(y)dy, so that pdF/dx=M(x) and pdF/pdy=N(y). So so by definition the equation is exact.

OpenStudy (deez):

i still dont get it.. can u type it as Equation formula please :)

OpenStudy (anonymous):

I just mean for a diff eqn to be exact it needs to be of the form Mdx+Ndy=0 with a function F(x,y) s.t. \[\delta F/\delta x = M\] and \[\delta F/\delta y = N\] (I can't find partial d). So if this case we already have M(x)dx+N(y)dy=0 so all we need is a function F. The function I wrote above works as this function, since the int Ndy disappears when partial diff by x and vice versa. So F works and the eqn is exact.

OpenStudy (deez):

Thanx alot .. helpful :)